OpenStudy (iwanttogotostanford):

HELP PLEASE IM STUCK: Using the completing-the-square method, find the vertex of the function f(x) = –2x^2 + 12x + 5 and indicate whether it is a minimum or a maximum and at what point. Maximum at (–3, 5) Minimum at (–3, 5)

jimthompson5910 (jim_thompson5910):

Vertex form is \(\Large y = a(x-h)^2+k\) The vertex of that equation is \(\Large (h,k)\) From \(\Large -2x^2 + 12x + 5\) we see that \(\Large a = -2, b = 12, c = 5\) Plug the values of 'a' and 'b' into the formula \[\Large h = \frac{-b}{2a}\] to find the value of h. Tell me what you get

OpenStudy (anonymous):

Oh, I thought you already know the answer to this. Hmmm.

OpenStudy (iwanttogotostanford):

so it would be A? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

why maximum?

OpenStudy (iwanttogotostanford):

because it increases

jimthompson5910 (jim_thompson5910):

you're looking at the graph I'm guessing?

OpenStudy (iwanttogotostanford):

yes

jimthompson5910 (jim_thompson5910):

do you know how to determine without the graph? by just looking at the value of 'a'

OpenStudy (iwanttogotostanford):

no, but can you just determine by looking at the graph?

jimthompson5910 (jim_thompson5910):

well let's say we don't have a graphing calculator since 'a' is a negative number, this means that the parabola opens downward like this  a good way to remember this is to think "The value of 'a' is negative. Negative means sad, so we have a sad face graph"

OpenStudy (iwanttogotostanford):

ok

jimthompson5910 (jim_thompson5910):

and of course, at the peak of this "sad face graph" is the vertex. Which in this case, is the max